Graph the function f(x)= x+1 if the domain is only negative integers?
Answer by glabow(165)

Consider what the function is if you let x=0. It graphs to the points on the line at y=1. (I'll assume you accept that y=f(x)=x+1.) But you can only allow x to take on values of less than 0 (the negative numbers), so the points of y will be all the points below the line for y=1. The range is the area below the line y=1, but not on it. And certainly not above it! To get there, x would have to be a positive number.
Now, you are only allowing negative integers in the domain. That means that the value of y will also be an integer. So, only the points below the line y=1 that have values of y that are integers are in the range. That would be the lines y=0, y=-1, y=-2, etc.
Done.